OpenStudy (anonymous):

y varies directly as x and inversely as the square of z. y=15 when x=80 and z=4. find y when x=81 and z=3

OpenStudy (unklerhaukus):

\[y(x,z)\propto xz^{-2}\]\[y(x,z)=kxz^{-2}\]

OpenStudy (unklerhaukus):

\[y(80,4)=k(80)(4)^{-2}=15\]

OpenStudy (unklerhaukus):

\[k=15\times4^2\div80\]

OpenStudy (unklerhaukus):

\[y(81,3)=(15\times4^2\div80)(81)(3)^{-2}=\cdots\]

OpenStudy (anonymous):

i came up with 3

OpenStudy (unklerhaukus):

did you understand the method /

OpenStudy (anonymous):

getting k first

OpenStudy (unklerhaukus):

yeah read 'varies directly' as 'proportional to' as 'is a constant multiple of'

OpenStudy (unklerhaukus):

three is good

OpenStudy (anonymous):

i dont think i am doing the equation right

OpenStudy (anonymous):

i also get 27 O_o

OpenStudy (unklerhaukus):

\[3^{-2}=\frac 1{3^2}\]

OpenStudy (anonymous):

i did 15*16/80

OpenStudy (anonymous):

=3

OpenStudy (unklerhaukus):

that is right

OpenStudy (anonymous):

thanks for the help

OpenStudy (unklerhaukus):

anytime